A major French insurer is in a position to take over an Irish-based insurance company called Farm Insure Ltd. Farm Insure has a particular focus on agricultural insurance which may complement the French insurer’s expertise in this sector. Using some preliminary information on Farm Insure, they can consider whether to continue their investigation into the financial position and examine whether the Irish insurer is a suitable take-over prospect.

1. With total assets of €242 million and total liabilities of €275 million, what is Farm Insure’s policyholders’ surplus?

2. Last year, Farm Insure had earned premiums of €77.5 million, but total revenues were €82.5 million Explain how this was possible. Why might this be the case?

3. Farm Insure Plus is the company’s main insurance product and it is sold to 35,000 farmers in Ireland. Farm Insure has incurred losses and loss adjustment expenses of €1,750,000 for this product over a 1-year period. What is the pure premium for Farm Insure Plus?

4. If the gross rate for Farm Insure Plus is €78.12, what is the expense ratio?

5. Although Farm Insure made a net underwriting loss last year, it paid €1.25 million in taxes. Explain why this was the case.
Answer #1

1. Policyholder surplus are assets of a policyholder-owned insurance company (also called a mutual insurance company) minus its liabilities. Policyholder surplus is one indicator of an insurance company’s financial health. Here we have less assets than liabilities thus there is no surplus

2. The interest earned by the investment of assets in either the general account for life insurance and property and casualty insurance companies or separate account for the life insurer is a component of overall revenue for the insurer. Thus remaining amount could be insurance

3. In the pure premium method, the pure premium is 1st calculated by summing the losses and loss-adjusted expenses over a given period, and dividing that by the number of exposure units. Thus pure premium would be €1,750,000/35000 = 50

4. Gross rate = pure premium + load. Thus load = 78.12-50 = 28.12. Expense ratio = load / gross rate = 28.12/78.12 = 36%